### v1

# Fluid flow inside slit-shaped nanopores: the role of surface morphology at the molecular scale

Abstract
Non-equilibrium molecular dynamics (NEMD) simulations of fluid flow have highlighted the peculiarities of nanoscale flows compared to classical fluid mechanics. In particular, boundary conditions can deviate from the no-slip behavior at macroscopic scales due to various factors. In this context, we investigate the influence of surface morphology in slit-shaped nanopores on the fluid flow. We demonstrate that the surface morphology effectively controls the slip length, which approaches zero when the molecular structures of the pore wall and the fluid are matched. Using boundary-driven, energy-conserving NEMD simulations with a pump-like driving mechanism, we examine two types of pore walls—mimicking a crystalline and an amorphous material—that exhibit markedly different surface resistances to flow. The resulting flow velocity profiles are consistent with Hagen–Poiseuille theory for incompressible, Newtonian fluids when adjusted for surface slip. For the two pores, we observe partial slip and no-slip behavior, respectively, which correlate with fluid layering and depletion near the surfaces. However, the confinement of the fluid gives rise to an effective viscosity that varies substantially with the pore width. Analysis of the hydrodynamic permeability shows that the simulated flows are in the Darcy regime. Additionally, the thermal isolation of the flow causes a linear increase in fluid temperature along the flow, which we relate to strong viscous dissipation and heat convection, utilizing conservation laws of fluid mechanics. Our findings underscore the need for molecular-scale modeling to accurately capture the fluid dynamics near boundaries and in nanoporous materials, where macroscopic models may not be applicable.
nanofluidics, capillary flow, hydrodynamic permeability, fluid state theory, non-equilibrium molecular dynamics simulation
I Introduction
In the past 60 years, molecular dynamics (MD) simulations have developed into an indispensable tool in many areas of research, complementing experimental investigations and theoretical predictions; examples range from drug development [1] and cell biology [2] to materials research [3, 4], chemical process engineering [5, 6], and geology [7]. Nowadays, it is used routinely to study the physical and chemical properties of complex materials and macromolecules as well as kinetic processes therein [8, 9, 10, 11]. This provides detailed insights into phase diagrams, structural properties, transformation rates, and transport coefficients, including dynamic shear viscosity, thermal and electrical conductivity, and diffusivity. The pioneering works [12, 13] employed classical particle models with simplistic interaction potentials to test fundamental predictions of statistical mechanics and kinetic theory. In this respect, the results of MD simulations have set and are still setting corner stones in our understanding of material responses, which were supported by and are still challenging the theoretical developments, a prime example being the long-time tails in the autocorrelation of particles in liquids [14, 15, 16, 17] and arrested fluids [18, 19, 20]. Moreover, the development of advanced, realistic force fields [21, 22, 23, 24] has turned MD simulations into a prime method for quantitative predictions of material properties, and these developments have rapidly gained new momentum with the advent of machine-learned interaction potentials that are based on first-principles calculations [25, 26, 27].
Whereas the investigated system sizes were originally limited to a few hundred atoms and just several thousand integration time steps, simulations of some million atoms over tens of millions of time steps can be performed routinely on today’s high-performance computing installations [28, 29, 30, 31], also owing to the massive parallelism provided by general-purpose accelerator hardware (“graphics processors”). This has brought small-wave number and low-frequency phenomena into the reach of MD simulations [32, 33, 16, 34, 28], thereby providing access to hydrodynamic scales with molecular resolution. It allows one to study set-ups typical of fluid mechanics in great detail such that also the boundary layers of a fluid flow can be resolved down to the scale of atomic interactions. This is particularly relevant for the development of realistic continuum models of pores via up-scaling approaches, which need information about the molecular behaviour at the nanoscale [35, 36]; there, hydrodynamic permeability models are prone to break down [37].
Transport phenomena in confined geometries, particularly at the nanoscale, showcase a continuously growing number of technological applications and relevance in natural processes [38, 39, 40]. The behavior of fluids in confined spaces is studied with the aim to design materials for filtration [41, 42], energy applications [37, 42, 43, 44], water desalination [45], chip-level cooling [46], and DNA sequencing [47], to name only a few. A critical factor in determining fluid behavior within a porous medium is the nature of the interactions between the fluid and the pore surfaces [48, 49]. Hydrophobicity, the tendency of a surface to repel water or similar polar fluids, modulates the fluid flow and distribution within the pore space; in particular, it can lead to pronounced slippage of the flow at the fluid–surface interface. Such slip, however, depends in a subtly way on factors such as surface roughness, wettability, and the presence of gaseous layers [50]. Surface roughness can change the fluid behavior significantly, which is known as the Lotus effect [49]: experiments evidence that specific surface coatings—such as polymers or nanotube structures, referred to as “nanoforests”—can create superhydrophobic surfaces with extremely low friction to flow [51]. The effect has also been demonstrated in MD simulations of fluid flow near nano-structured surfaces [52]. On the other hand, for hydrophilic surfaces with a similar topography, one may expect that trapping of fluids within surface asperities or pockets creates additional friction and reduces slippage.
For atomically smooth, crystalline surfaces, it was shown that hydro- and solvophobicity facilitate fluid slip [53], which was linked to the formation of distinct interfacial layers, where fluid density and structure differ significantly from those in the bulk [54]. However, despite a general trend that hydrophobic surfaces show a reduced fluid–surface interaction and enhance fluid slip, hydrophobicity alone does not imply a large slip length and vice versa [55]. Moreover, subtle changes in the fluid–surface interaction can drive significant structural and density changes near the surface without modifying the wetting properties of the surface. There is no simple correlation between the interfacial fluid structure and the hydrophobicity of the surface and we refer to the partially controversial discussion in, e.g., Refs. 56, 57, 58. Also atomic-level details like surface composition, molecular orientation, and electronic interactions can profoundly impact the slip behavior, even on structurally similar surfaces, which was demonstrated in ab initio simulations of water near graphene surfaces [59, 60].
Fluid flow through nanoporous media using non-equilibrium MD (NEMD) simulations has been investigated with low intensity since the 1990’s (e.g., Ref. 61), but has seen a surge in research activity during the past decade. The focus is typically on the permeability of ordered porous media such as MOFs [62, 63] and linear or slit-shaped pores [54, 64, 65, 66, 67, 68], but also through disordered organic materials [69, 70, 37]. The majority of these studies addresses the transport of molecules in the gas phase or of low-density fluids, while NEMD simulations for dense liquids face particular technical challenges. One approach to steady fluid flow is to prescribe the chemical potential of the inlet and outlet by means of grand-canonical (GC) reservoirs (dual control volume set-up) [71, 69, 70]. However, the Monte Carlo part of this approach would become inefficient for dense liquids due to the low acceptance rate of trial insertions of particles. One solution to this problem is to control the pressure instead of the chemical potential [72], and a recent, reservoir-based alternative to GC-NEMD simulations is offered by spatially adaptive resolution simulations (AdResS) [73, 74, 75]. Another line of approach to steady flow is driving the liquid in a boundary region of a periodic set-up by applying an external force to maintain a pressure difference [67, 68] or a concentration gradient [63]. More generally, custom flow fields can be generated by extending the force field into the whole pore space [76]; in the context of incompressible flows, a constant force was used to mimic a constant pressure gradient [54, 65]. Below, we describe boundary-driven NEMD simulations of steady flow implement via a pump-like driving mechanism, which is applied to the particles in a small region at the entrance of the pore, while preserving the Hamiltonian character of the fluid dynamics in the remaining pore volume.
No-slip boundary conditions are a customary and realistic choice in fluid engineering to model fluid flow through a macroscopic conduit. On the other hand, the above discussion shows that the control of specific boundary conditions of fluid flow in MD simulations is a challenging task. The ability to realize no-slip boundaries at the molecular scale would facilitate the comparison of particle-resolved simulations with predictions from continuum fluid dynamics. In particular, one may expect that the latter predictions apply to a larger part of the simulated volume, which can serve as a coarse-grained particle model of fluid flow with low compressibility. A possible route to this aim are surfaces with a molecular roughness in order to introduce tangential stresses on the fluid, and thereby increasing the friction with the surface and decreasing the slip.
In this work, we investigate how surface morphology, or molecular-scale surface topography, affects fluid flow near the surface. In particular, how does it modify the effective boundary conditions, the local structure of the fluid, and the overall permeability? We propose and characterize a nanopore model with discrete walls composed of obstacles arranged to match the molecular structure of the fluid. We show that such a configuration allows efficient control of surface slip, which is almost completely eliminated. More specifically, we simulate the steady flow of a simple fluid inside a slit-shaped nanopore as a minimal model of a porous medium; this choice enables comparison with previous studies and has obvious extensions to other pore cross sections. We relate the molecular-level observations of fluid flow to the basic laws of fluid mechanics and test predictions of Hagen-Poiseuille theory for incompressible, Newtonian fluids. Lastly, we discuss the issue of heat transport in isolated fluid flow under conditions of strong viscous dissipation, as encountered in fast-flow scenarios.
II Fluid mechanical considerations
II.1 Hagen–Poiseuille flow in a slit-shaped conduit
We consider the laminar flow of a fluid through a slit-like conduit of width $D$ and square-shaped surfaces, each of macroscopic area $L^{2}$. The application of a pressure difference $\Delta P=P_{\text{in}}-P_{\text{out}}$ at the inlet and outlet of the conduit leads to a Hagen–Poiseuille flow, which we shall briefly re-derive here. To this end, we choose a coordinate frame such that the $x$-axis points parallel to the wall surfaces, in the direction of the flow, and the $z$-axis is perpendicular to the surfaces, which are located at $z=\pm D/2$ so that $z=0$ denotes the plane in the center of the flow. By translational symmetry, the problem does not depend on the $y$-coordinate of the position.
The steady flow of a uniform, incompressible, Newtonian fluid is governed by the stationary Navier–Stokes equation [77, 78]:
$m\varrho\left({\boldsymbol{v}}\cdot\nabla\right){\boldsymbol{v}}=-\nabla P+\eta\nabla^{2}{\boldsymbol{v}},$ (1)
with $m\varrho(x,z)$ being the fluid density field, ${\boldsymbol{v}}(x,z)$ the velocity field, $P(x,z)$ the pressure field, and $\eta$ the dynamic shear viscosity. At macroscopic scales, no-slip boundary conditions hold at the surfaces of the conduit:
${\boldsymbol{v}}(x,\pm D/2)=0.$ (2)
At low Reynolds number, the l.h.s. of Eq. 1 is negligible and the velocity field is aligned with the direction of the mean flow, ${\boldsymbol{v}}=(v_{x},0,0)$. Due to the incompressibility of the fluid, low Mach number, and in the absence of an external force, the pressure gradient is constant: $\partial_{x}P(x,z)=-\Delta P/L$. Thus, the problem becomes $x$-independent and the $x$-component of the flow field obeys:
$0=\Delta P/L+\eta\,\partial^{2}_{z}v_{x}(z)\,.$ (3)
Applying the boundary condition in Eq. 2 results in the Hagen–Poiseuille flow profile [79, 80]:
$v_{x}(z)=\frac{\Delta P}{2\eta L}\left(\frac{1}{4}D^{2}-z^{2}\right).$ (4)
Unlike in macroscale systems, where the friction near surfaces is fully developed, at the atomistic scale, the fluid molecules slip along the surface and, the arrangement of surface atoms, as well as their interaction with the fluid molecules, affects the boundary layer of the flow. These lead to an effective full or partial slip of the flow field, described by Navier’s partial slip condition [77, 81]:
$v_{x}(z=\pm D/2)=\mp\beta_{\text{slip}}\,\partial_{z}v_{x}(z=\pm D/2),$ (5)
where $\beta_{\text{slip}}$ is the slip coefficient, representing the distance below the surface at which the no-slip condition would hold if the velocity was linearly extrapolated into the wall. As an alternative to $\beta_{\text{slip}}$, we introduce the slip length ${\delta_{\text{slip}}}$ as an apparent displacement of the pore surface from, e.g., $z=D/2$ to $z=D/2+{\delta_{\text{slip}}}$. The velocity profile for partial slip follows from solving Eq. 3 for the boundary condition (5) instead of the no-slip condition Eq. 2:
$v_{x}(z)=\frac{\Delta P}{2\eta L}\left(\frac{1}{4}(D+2{\delta_{\text{slip}}})^{2}-z^{2}\right);$ (6)
in this case, the two definitions of the slip length are related via:
$\beta_{\text{slip}}={\delta_{\text{slip}}}(1+{\delta_{\text{slip}}}/D)\,.$ (7)
In many applications, interest often lies in the total amount of fluid passing through the conduit per unit time, which is quantified by the volumetric flow rate or, equivalently, the mass flux per unit area,
$J_{\varrho}=\frac{1}{D}\int_{-D/2}^{D/2}\varrho v_{x}(z)\,\mathrm{d}z\,,$ (8)
which is given as the momentum density field, $\varrho{\boldsymbol{v}}$, averaged over the cross section of the conduit. In the case of Hagen–Poiseuille flow without slip, one finds:
$J_{\varrho}=\frac{2}{3}\varrho v_{x}(0)=\frac{D^{2}\Delta P}{12\nu L}\,,$ (9)
using Eq. 4 and the kinematic viscosity $\nu=\eta/\varrho$. In the field of porous media, the proportionality between fluid flow and the applied pressure gradient is known as Darcy’s law[82, 48], which defines the permeability $k$ of the medium:
$J_{\varrho}=\frac{k\Delta P}{\nu L}\,.$ (10)
Comparison with Eq. 9 yields the permeability of the slit-shaped conduit:
$k=\frac{D^{2}}{12}.$ (11)
If slip at the surface is taken into account, one calculates from Eq. 6:
$k$ $=\frac{D^{2}}{12}+\beta_{\text{slip}}(D/2)$ (12)
$=\frac{1}{12}\left(D^{2}+6D{\delta_{\text{slip}}}+6{\delta_{\text{slip}}}^{2}\right),$ (13)
showing that the presence of slip increases the effective permeability of the pore, allowing more fluid to pass through for the same pressure gradient.
II.2 Non-isothermal flow
For the interpretation of data obtained within energy-conserving MD simulations, as presented below, we need to relate viscous dissipation and heat conduction to the fluid temperature. Such a relation can be found in fluid mechanics texts [78, 83]; here, we will only sketch its derivation. To this end, we start from the first law of thermodynamics applied to a small control volume $V$. The total energy $E$ within the volume consists of the internal energy due to molecular interactions and thermal motion, and the kinetic energy due to the macroscopic fluid flow, but it does not include the potential energy due to external forces. According to the first law, the rate of change $\mathrm{d}E/\mathrm{d}t$ of the energy balances the rate $\dot{Q}$ of heat added to the volume minus the rate $\dot{W}$ of work done by the volume:
$\mathrm{d}E/\mathrm{d}t=\dot{Q}-\dot{W}\,.$ (14)
The energy content of the control volume changes due to local energy conversion and due to advection with the flow field:
$\frac{dE}{dt}=\int_{V}\partial_{t}\bigl{(}m\varrho e+\tfrac{1}{2}m\varrho|{\boldsymbol{v}}|^{2}\bigr{)}\,\mathrm{d}V\\ +\int_{\partial V}\bigl{(}m\varrho e+\tfrac{1}{2}m\varrho|{\boldsymbol{v}}|^{2}\bigr{)}{\boldsymbol{v}}\cdot{\boldsymbol{n}}\,\mathrm{d}A,$ (15)
where the field $e$ denotes the internal energy per unit mass, $m\varrho$ is the mass density field, and ${\boldsymbol{v}}$ the flow field, as above. The work performed by the volume has a contribution stemming from the pressure acting on its surface $\partial V$ and a contribution due to viscous forces at the surface. For the rate of work:
$\dot{W}=\int_{\partial V}P{\boldsymbol{v}}\cdot{\boldsymbol{n}}\,\mathrm{d}A-\int_{\partial V}{\boldsymbol{v}}\cdot(\boldsymbol{\tau}\cdot{\boldsymbol{n}})\,\mathrm{d}A\,,$ (16)
where the unit vector ${\boldsymbol{n}}$ is the outward normal to the surface and the symmetric tensor $\boldsymbol{\tau}$ is the deviatoric stress of the fluid. Moving from the rate of heat, $\dot{Q}$, to that of the heat density, per volume, $\dot{q}$, and transforming the surface integrals into volume integrals via Gauss’ divergence theorem yields a local form of energy conservation:
$\partial_{t}\bigl{(}m\varrho e+\tfrac{1}{2}m\varrho|{\boldsymbol{v}}|^{2}\bigr{)}+\nabla\cdot\bigr{[}m\varrho{\boldsymbol{v}}\bigl{(}e+P/(m\varrho)+\tfrac{1}{2}|{\boldsymbol{v}}|^{2}\bigl{)}\bigl{]}\\ =\dot{q}+\nabla\cdot({\boldsymbol{v}}\cdot\boldsymbol{\tau}).$ (17)
The l.h.s. is further simplified by expanding the derivatives over the products and introducing the material derivative, $\mathrm{D}_{t}=\partial_{t}+{\boldsymbol{v}}\cdot\nabla$, as well as by making use of the continuity equation for mass conservation,
$\partial_{t}(m\varrho)+\nabla\cdot(m\varrho{\boldsymbol{v}})=\mathrm{D}_{t}m\varrho+m\varrho\nabla\cdot{\boldsymbol{v}}=0\,$ (18)
and Cauchy’s momentum equation, which yields:
$m\varrho\mathrm{D}_{t}{\boldsymbol{v}}=-\nabla P+\nabla\cdot\boldsymbol{\tau}\,.$ (19)
On the r.h.s., we rewrite $\nabla\cdot({\boldsymbol{v}}\cdot\boldsymbol{\tau})={\boldsymbol{v}}\cdot(\nabla\cdot\boldsymbol{\tau})+\Phi$ upon defining the viscous dissipation function:
$\Phi=\boldsymbol{\tau}:\nabla{\boldsymbol{v}}=\tau_{\alpha\beta}(\partial_{\alpha}v_{\beta})\,;$ (20)
the sum convention is applied in the last expression. Collecting terms leads to the energy equation, which holds generally, here, in the absence of external forces, since it is merely an expression of the fundamental conservation laws:
$m\varrho\mathrm{D}_{t}e+P\nabla\cdot{\boldsymbol{v}}=\dot{q}+\Phi\,.$ (21)
Assuming thermodynamic equilibrium locally, in each control volume, spatial changes of the internal energy $e$ are determined by the fluid’s caloric and thermal equations of state, $e=e(T,\varrho)$ and $P=P(T,\varrho)$. We invoke the fundamental relation [84, 85]:
$\mathrm{d}e(T,\varrho)=c_{v}(T,\varrho)\,\mathrm{d}T-P(T,\varrho)\,\mathrm{d}(\varrho^{-1})\,,$ (22)
where the irreversibly added heat $\delta q=c_{v}\mathrm{d}T$ is expressed through the isochoric specific heat $c_{v}=(\partial e/\partial T)_{\varrho}$ and $\varrho^{-1}$ denotes the specific volume. Then, the chain rule for $\mathrm{D}_{t}$ implies:
$\mathrm{D}_{t}e(T,m\varrho)$ $=\left(\frac{\partial e}{\partial T}\right)_{\!\varrho}\mathrm{D}_{t}T+\left(\frac{\partial e}{\partial\varrho}\right)_{\!T}\mathrm{D}_{t}m\varrho$
$=c_{v}\mathrm{D}_{t}T+m\varrho^{-2}P\,\mathrm{D}_{t}\varrho\,.$ (23)
Thus, the l.h.s. of Eq. 21 can be rewritten to yield the temperature form of the energy equation:
$c_{v}m\varrho\mathrm{D}_{t}T=\dot{q}+\Phi\,,$ (24)
where we have used mass conservation, Eq. 18, to cancel the term $P\nabla\cdot{\boldsymbol{v}}$. Equation 24 rests on mild assumptions and, in particular, holds for non-uniform, compressible, non-Newtonian fluids. It remains to specify the terms on the r.h.s. of the equation.
[…]
${\boldsymbol{j}}_{q}=-\kappa\nabla T,$ (25)
which approximates the heat flux ${\boldsymbol{j}}_{q}$ as the linear response to a (small) temperature gradient $\nabla T$, in accordance with experimental observations for both fluids and solids; the coefficient $\kappa$ is the thermal conductivity. Hence,
$\dot{q}=\kappa\nabla^{2}T.$ (26)
For the viscous dissipation function, Eq. 20, we specialise to an incompressible, Newtonian fluid. Under these conditions the viscous stress tensor is $\boldsymbol{\tau}=\eta\dot{\boldsymbol{\gamma}}$, where the rate of strain tensor is given by $\dot{\boldsymbol{\gamma}}=\nabla{\boldsymbol{v}}+(\nabla{\boldsymbol{v}})^{\top}$. Accordingly, the viscous dissipation function reads:
$\Phi$ $=\eta(\partial_{\alpha}v_{\beta}+\partial_{\beta}v_{\alpha})(\partial_{\alpha}v_{\beta})=\tfrac{1}{2}\eta\dot{\boldsymbol{\gamma}}:\dot{\boldsymbol{\gamma}}\,,$ (27)
and substituting into Eq. 24 yields to the equation of the temperature field under the named approximations [83]:
$c_{v}m\varrho(\partial_{t}T+{\boldsymbol{v}}\cdot\nabla T)=\kappa\nabla^{2}T+\tfrac{1}{2}\eta\dot{\boldsymbol{\gamma}}:\dot{\boldsymbol{\gamma}}\,.$ (28)
It describes how the local temperature of the fluid changes as a result of fluid convection (${\boldsymbol{v}}\cdot\nabla T$), heat conduction within the fluid ($\kappa\nabla^{2}T$), and viscous dissipation ($\eta\dot{\boldsymbol{\gamma}}:\dot{\boldsymbol{\gamma}}/2$); the latter accounts for the irreversible work performed due to friction between fluid layers.
III Simulation Method
This study employs boundary-driven non-equilibrium molecular dynamics (NEMD) simulations to explore nanoscale mass transport of a simple fluid within slit-shaped pore structures formed by discrete obstacles. The simulations were conducted using HAL’s MD package [88, 89], a software that leverages the high parallelism of graphics processors (GPUs) and other accelerator hardware [90] and that is an efficient and precise tool for studying inhomogeneous fluids [32, 91, 34, 92, 93].
III.1 Fluid model
For the fluid, we employ a simple liquid consisting of repulsive spheres interacting via a smoothed Weeks–Chandler–Andersen (WCA) potential:
$U_{\text{WCA}}(r)=\left[U_{\text{LJ}}(r)-U_{\text{LJ}}(r_{c})\right]\,g\boldsymbol{(}(r-r_{c})/h\boldsymbol{)},$ (29)
which is derived from the Lennard-Jones pair potential, $U_{\text{LJ}}(r)=4\varepsilon\bigl{[}(\sigma/r)^{12}-(\sigma/r)^{6}\bigr{]}.$ The parameters $\varepsilon$ and $\sigma$ serve as units of energy and length, respectively, and the unit of time is chosen to be $\tau=\sqrt{m\sigma^{2}/\varepsilon}$ in terms of the particle mass $m$. Placing the cutoff distance $r_{c}$ at the minimum of the potential, $r_{c}=2^{1/6}\sigma$, the interaction contains only the repulsive forces that prevent particle overlap without considering attractive dispersion forces. A smooth truncation is used to ensure good energy conservation during long MD simulations in the microcanonical (NVE) ensemble, implemented via multiplication of the energy-shifted potential with the algebraic interpolation function [94, 95, 89, 33] $g(x)=x^{4}/\bigl{(}1+x^{4}\bigr{)}$ for $x\leqslant 0$ and $g(x)=0$ otherwise. Thereby, $U_{\text{WCA}}(r)$ is rendered $C^{2}$-continuous also at the cutoff. We choose $h=0.005\sigma$ for the truncation range, which is small enough to entail only minor changes of the fluid properties due to the truncation. Throughout this study, density and temperature of the quiescent fluid, without driving, are fixed to $\varrho=0.7\sigma^{-3}$ and $T=1\varepsilon/k_{\text{B}}$.
III.2 Pore morphologies
Figure 1: Top: sketch of the geometry of the simulated slit-shaped pore space. The cuboid is filled with fluid, which is driven at the pore entrance (yellow region) in the direction of the thick arrow, indicating the flow. The flow is delimited by the pore walls at the top and bottom faces of the cuboid, and periodic boundary conditions are applied at the other faces.  Bottom: Snapshots of the particle configurations in NEMD simulations for two different obstacle arrangements (red particles), which realize walls with a crystalline (fcc) and amorphous structure, respectively. The pore is filled with fluid particles (blue). For better visibility, a smaller pore space ($L=50$) than the one used in the actual simulations is shown.
[…]
The pore walls are implemented as arrangements of obstacle particles, which are fixed in space and which exert forces on the fluid particles according to the same WCA potential, Eq. 29, that is used for the interaction between fluid particles. The use of such discrete walls provides us with flexibility regarding the surface morphology. This allows us to study the role of morphology in the fluid flow by comparing simulations with two types of pore surfaces, which we refer to as crystalline and as amorphous walls.
For crystalline walls, the pore particles are placed on a face-centered cubic (fcc) lattice with a lattice constant of $a_{\text{lat}}=1.25\sigma$, Fig. 1. (Technically, the dimension of the simulation box is extended along the $z$-axis by adding $15\sigma$ in total to accommodate the obstacle particles.) The lattice constant is chosen small enough to prevent infiltration of fluid particles into the wall structure, it implies nearest-neighbor distances of $a_{\text{lat}}/\sqrt{2}\approx 0.88\sigma$. The dense arrangement of pore particles results in a highly ordered, crystalline structure, exposing a comparably smooth surface to the fluid.
For amorphous walls, the pore particles are distributed in a disordered fashion in such a way that the fluid structure is matched, Fig. 1. Such structures are generated by filling the whole simulation volume, $(D+15\sigma)\times L\times L$, with a WCA fluid at the desired fluid density and temperature and equilibrating its structure in a short NVT simulation run. Afterwards, the particles found outside the (nominal) volume of the conduit are identified as obstacles and they delimit the pore space, while those inside the pore region represent the fluid; obstacle particles do not participate in the further integration of the equations of motion. The procedure leads to comparably rough, irregular wall surfaces in contact with the fluid.
The presence of density fluctuations in the obstacle configuration and the overall lower obstacle density, we applied an additional measure to prevent fluid particles from penetrating the walls. Alongside the obstacles, two flat, planar walls are positioned at a distance of $\sigma$ behind the pore surfaces. For the interaction energy of a fluid particle at position $z$ with each wall, we use a purely repulsive surface potential:
$U_{\text{flat}}(\hat{z}<r_{c})=\frac{4\pi\varepsilon_{w}}{3\times 15}\Bigl{(}\frac{\sigma}{\hat{z}}\Bigr{)}^{9},$ (30)
in terms of the distances $\hat{z}=z-(D/2+\sigma)$ and $\hat{z}=z+(D/2+\sigma)$. The surface potential is smoothly truncated, analogously to Eq. 29, using again the cutoff distance $r_{c}=2^{1/6}\sigma$ with $h=0.005\sigma$.
III.3 Boundary-driven NEMD simulation
Thermostatted pump.
Boundary-driven flows are generated in the NEMD simulations by placing a “pump”, combined with a thermostat, at the entrance of the pore space [75]. For particles in the region $0\leqslant x\leqslant L_{\text{pump}}=10\sigma$ along the $x$-axis, new velocities are randomly assigned, after fixed intervals of $\tau_{\text{pump}}=0.2\tau$, by sampling from a shifted Gaussian distribution with mean $v_{0}{\boldsymbol{e}}_{x}$ and variance $k_{\text{B}}T/m$; the latter corresponds to the Maxwell–Boltzmann distribution at temperature $T$ with an imposed flow velocity $v_{0}$ along the $x$-axis. Thereby, particles leave the pump region with the desired target velocity, pushing away particles in front of them and thereby creating a steady mass flux in the pore. We note that due to the resistance of the pore to the flow, the achieved average flow velocity is slightly below the nominal pump velocity $v_{0}$ (see below). This issue can be mitigated by decreasing the update interval $\tau_{\text{pump}}$. For the data interpretation, we will work with the achieved flow velocity.
The motion of all fluid particles, including those in the pump region, is generated with the velocity Verlet integrator, yielding a Hamiltonian dynamics in the pore space ($L_{\text{pump}}\leqslant x\leqslant L$), except in the pump region; in particular, mass, momentum, and energy are locally conserved.
Simulation protocol.
The system preparation starts with an equilibration phase in a canonical (NVT) ensemble over a duration of $100\tau$, where the number of particles and the volume of the simulation box are fixed and a global Andersen thermostat [96], with an update rate of $10\tau^{-1}$, is used to control the temperature. After equilibration, the simulation transitions to the non-equilibrium set-up, by replacing the Andersen thermostat with the thermostatted pump in the boundary region. Two NEMD simulation runs follow, each of a duration of $100\tau$, first to relax the fluid to the non-equilibrium steady state and, finally, to collect the data for the observables. For equilibrium MD simulations, we have chosen an integration timestep of $\delta t=0.001\tau$, which was increased to $\delta t=0.005\tau$ in NEMD simulation runs.
III.4 Observables
Spatially resolved properties.
During data collection, phase space trajectories, combining particle positions and velocities, are saved regularly after intervals of length $\tau$ to the disk using the H5MD file format [97]. Spatially resolved profiles of the flow velocity are calculated along the direction normal to the pore surfaces. To this end, the pore space is divided along the $z$-axis into equally-spaced slabs of width $0.375\sigma$ and centered at the positions $z_{n}$. For each of these “bins”, the set $S_{n}(t)$ of particles located in bin $n$ at time $t$ is determined and the mean velocity $v_{x}(z_{n})$ is computed from:
$v_{x}(z_{n})=\frac{1}{\langle|S_{n}(t)|\rangle}\Bigl{\langle}\sum_{j\in S_{n}(t)}u_{x,j}(t)\Bigr{\rangle},$ (31)
where $|S_{n}(t)|$ represents the number of particles in the set $S_{n}(t)$, and $u_{x,j}(t)$ is the $x$-component of the velocity of particle $j$ at time $t$.
[…]
Figure 2: Exemplary transversal profiles of the fluid velocity $v_{x}(z)$ at different $x$ positions along the flow direction. The lines show parabolic fits to simulation data obtained for a wide slit-shaped pore ($D=60\sigma$) with amorphous walls. A mean velocity of $v_{0}=0.5\sigma\tau^{-1}$ is imposed on the fluid in the pump region, at the entrance of the pore ($x<10\sigma$). One observes that a plug-like flow at the entrance develops into the classical parabolic shape of Hagen–Poiseuille flow downstream.
The fluid particles leave the thermostatted pump region with a spatially homogeneous mean velocity, which does not depend on the distances to the pore surfaces. This corresponds to the plug-like flow that is mathematically obtained for (almost) full-slip boundary conditions at the surfaces. Only gradually, the plug flow transitions downstream into Hagen–Poiseuille flow due to viscous stresses. This transformation is exemplified in Fig. 2 for the flow inside a wide pore ($D=60\sigma$) and a pumping velocity of $v_{0}=0.5\sigma\tau^{-1}$. The transversal velocity profiles $v_{x}(z)$ at different $x$ positions exhibits an almost flat behavior near $x=L_{\text{pump}}$, with little velocity variation, which changes into a parabolic shape for $x\gtrsim 30\sigma$ as the flow develops towards the center of the pore space. This issue is less relevant for more narrow pores, i.e., the Hagen–Poiseuille flow is already well developed closer to the pump region. In the computation of quantities averaged along the flow, we take this effect into account by discarding the upstream part of the pore space and considering only particles with $x>50\sigma$.
Density profile via Fourier modes.
For the spatial variation of the density perpendicular to the wall surfaces, we used a different route based on a Fourier representation of the fluctuating density field, encoding all particle positions: $\hat{\varrho}(z)=\sum_{j=1}^{N}\delta(z-z_{j})$; here, it suffices to consider only the Cartesian component along the $z$-axis. First, the (fluctuating) density modes [98]:
$\hat{\varrho}_{k_{z}}(t)=\sum_{j=1}^{N}\exp(ik_{z}z_{j}(t))\,$ (32)
are computed for all wave “vectors” $k_{z}$ in parallel, by summing the contributions from all $N$ particles, positioned at $z_{j}(t)$ at time $t$; these summations occur during the simulation and take full advantage of GPU acceleration. The wave vectors are located on the reciprocal lattice of the periodic domain of the simulation; here, $k_{z}=2\pi n/L_{z}$ for $|n|\leqslant n_{\text{max}}$ integer and $L_{z}$ is the length of the simulation box in the $z$-direction. The stationary density profile in real space, $\varrho(z)$, follows from applying the corresponding Fourier backtransform to the time-averaged density modes:
$\varrho(z)=L_{z}^{-1}\sum_{k_{z}}\langle\varrho_{k_{z}}(t)\rangle\,\text{e}^{-ik_{z}z-w^{2}k^{2}}\,,$ (33)
carried out once in a post-processing step. The spatial resolution $\Delta z$ of the method is determined by the maximum wave number $k_{\text{max}}=2\pi n_{\text{max}}/L_{z}$, which was chosen to be $n_{\text{max}}=120$ for $L_{z}=30\sigma$, corresponding to $\Delta z=\sigma/4$. Equation 33 includes a Gaussian filter of width $w=\Delta z/2$, which slightly smoothens the density profile $\varrho(z)$ by softly truncating the spectrum at long wave numbers, thereby avoiding spurious small-wave length oscillations.
Shear viscosity.
The shear viscosity $\eta_{b}$ of the bulk fluid was computed in equilibrium MD simulations [33] of the bulk fluid, which is homogeneous and isotropic, using the Einstein–Helfand formula [99, 100],
$\eta_{b}=\frac{\varrho}{k_{\text{B}}T}\lim_{t\to\infty}\frac{1}{6N}\partial_{t}\sum_{\alpha<\beta}\Bigl{\langle}\Bigl{(}\int_{0}^{t}\!\Pi_{\alpha\beta}(t^{\prime})\,\mathrm{d}t^{\prime}\Bigr{)}^{2}\Bigr{\rangle},$ (34)
for Helfand’s second moment of the time-integrated, microscopic expression for the total stress tensor [98],
$\boldsymbol{\Pi}(t)=\sum_{i=1}^{N}\Bigl{[}m{\boldsymbol{u}}_{i}(t)\otimes{\boldsymbol{u}}_{i}(t)+\sum_{j>i}^{N}{\boldsymbol{r}}_{ij}(t)\otimes{\boldsymbol{F}}_{ij}(t)\Bigr{]}\,,$ (35)
where ${\boldsymbol{u}}_{i}(t)$ is the velocity of particle $i$, ${\boldsymbol{r}}_{ij}(t)={\boldsymbol{r}}_{i}(t)-{\boldsymbol{r}}_{j}(t)$ the separation vector between particles $i$ and $j$ and, ${\boldsymbol{F}}_{ij}(t)$ the force acting on particle $j$ due to particle $i$, and $\otimes$ denotes the tensor product. For the repulsive fluid at density $\varrho=0.7\sigma^{-3}$ and temperature $T=1\varepsilon/k_{\text{B}}$ (Section III.1), we have obtained the bulk value of the shear viscosity to be:
$\eta_{b}=(1.417\pm 0.004)\varepsilon\tau\sigma^{-3}.$ (36)
With this, we can characterize the flow regime in terms of the Reynolds number:
$\text{Re}=\frac{m\varrho\bar{v}D}{\eta_{b}},$ (37)
[…]
IV Results & Discussion
IV.1 Surface layering in transversal density profiles
Figure 3: Number density $\varrho(z)$ of the simulated fluid flow as a function of the distance $\tilde{z}=\left(z+D/2\right)$ to one of the pore surfaces (vertical, dashed lines); the surfaces stem from (a) crystalline and (b) amorphous walls. Colors distinguish the results for six nominal pump velocities $v_{0}$, including the equilibrium case ($v_{0}=0$).
In a first step, we determined the implications of the morphology of the pore surfaces on the fluid structure near the surfaces. For the two kinds of surfaces, as realized by crystalline and amorphous walls (Section III.2), we have calculated the number density profiles $\varrho(z)$ as a function of the distance $\tilde{z}=z\pm D/2$ to the surfaces, Fig. 3, both in equilibrium ($v_{0}=0$) and for fluid flows driven at nominal velocities between $v_{0}=0.1\sigma\tau^{-1}$ and $1\sigma\tau^{-1}$. For the calculation, only the $z$-component of the particle positions was used, Eq. 32, and the profiles shown are thus averaged along the $x$- and $y$-axes, i.e., along the directions parallel to the surfaces.
A first observation is that the density profiles remain largely unaffected by changes in the flow velocity; even for the strongest driving studied, they correspond to the equilibrium profiles. This suggests, first, that the flowing fluid is locally in equilibrium and, second, that the obtained density profile can uniquely be associated to an effective surface potential [101, 102], which may have both attractive and repulsive components. In this context, we recall that the simulated pore walls are made of purely repulsive obstacles that are merely arranged differently for the two wall types.
In case of the densely packed, crystalline walls, $\varrho(z)$ displays three to four distinct monolayers of adsorbed fluid near the surfaces, which resemble the layering phenomena known for, e.g., hard spheres near a flat, impenetrable surface [103, 56]. (As a side remark, such layering is absent in the composition profile for surface adsorption of binary fluids [93].) We interpret the behavior of $\varrho(z)$ as a direct consequence of the geometric regularity and dense packing of the crystalline obstacle configuration, shown in simulation snapshot in Fig. 1. In contrast, the density profile lacks any pronounced peaks for the amorphous walls, Fig. 3b), with the obstacle configuration matching the fluid structure: $\varrho(z)$ remains almost constant upon approaching the surface as close as $\tilde{z}\approx 1\sigma$ and, finally, decays to zero monotonically. We conclude that the use of amorphous walls in simulations can essentially eliminate the issue of layering effects, so that bulk-like behavior can be expected in a larger volume of the confined fluid. The depletion zone observed at the amorphous wall is reminiscent of a gaseous layer at the surface, which may be related to solvophobicity, but not necessarily, and which would suggest a low resistance to fluid flow, i.e., a large slip length.
IV.2 Hagen–Poiseuille flow with slip
Figure 4: Variation of the flow velocity $v_{x}(z)$ transversal to flow, as function of the distance $z$ to the pore center. Symbols show NEMD simulation data for a slit pore of a nominal width of $D=15\sigma$ with (a) crystalline and (b) amorphous walls; for clarity, only every second item data points are shown. Colors distinguish the results for six nominal pump velocities $v_{0}$; the realized average flow velocities $\bar{v}_{x}$ are indicated by the bars on the left. Solid lines are parabolic fits of the Hagen–Poiseuille law for partial slip boundary conditions, Eq. 6, to the disk-shaped symbols; black circles represent data excluded from the fits.
Figure 5: (a) Slip length ${\delta_{\text{slip}}}$ and (b) effective shear viscosity $\eta$ obtained from the fits to the velocity profiles in Fig. 4 for Hagen–Poiseuille flow between crystalline, red diamonds, and amorphous, blue disks, pore walls. Solid lines connect the data points and serve as guide for the eye.
The surface morphology of the pore walls also determines the flow velocity $v_{x}(z)$ in the cross section of the pore, as a function of the distance to the pore center at $z=0$. For the crystalline walls, the velocity profiles of the simulated fluid flows exhibit a parabolic shape in the pore space for a range of pumping velocities $v_{0}$, Fig. 4a). However, close to the surfaces, the obtained flow velocities are still considerably large, approximately one half of the maximum value $v_{x}(0)$. The data points next to the surface positions display jumps to values close to zero, which we attribute to the very low particle densities $\varrho(z)$ at these positions and thus a poor statistics. For all pumping velocities, the velocity profiles $v_{x}(z)$ are well described by the modified Hagen–Poiseuille law, Eq. 6, which permits also a non-zero slip length ${\delta_{\text{slip}}}$. The values of ${\delta_{\text{slip}}}$ fitted to the data are on the order of two fluid particle diameters and depend on the strength of the fluid flow; they increase mildly from ${\delta_{\text{slip}}}\approx 1.23\sigma$ to $2.13\sigma$ upon increasing $v_{0}$ from $0.1\sigma\tau^{-1}$ to $1\sigma\tau^{-1}$, Fig. 5a). This pronounced deviation from the no-slip boundary condition, expected to hold at macroscopic scales, suggests an only moderate resistance of the pore walls to the fluid’s movement along the surfaces. One possible contribution to the enhancement of the surface’s flow is a putative sliding motion of the distinct layers that occur near the crystalline walls, Fig. 3a).
In contrast, the amorphous walls yield nearly perfect no-slip boundary conditions for the flow, as shown in Fig. 4b). The obtained slip lengths are an order of magnitude smaller than for the crystalline walls and increase from ${\delta_{\text{slip}}}\approx 0.17\sigma$ to $0.30\sigma$ upon increasing the flow velocity, Fig. 5a). Thus, fluid particles encounter a strong resistance as they flow past by the disordered surface, which we attribute to the disordered, fluid-like structure of this type of wall, generating an optimal roughness at the atomic scale.
IV.3 Effective viscosity of the confined fluids
Figure 6: Variation of fluid pressure $P(x)$, temperature $T(x)$, and number density $\varrho(x)$ along the direction of the flow inside a slit-shaped nanopore with either crystalline (top row) or amorphous (bottom row) walls. The pore is of width $D=15\sigma$ and has a length of $L=100\sigma$; periodic boundary conditions are applied along the $x$-axis. Disks represent simulation data and solid lines denote linear fits to the data for $x>50\sigma$, indicating constant pressure and temperature gradients, respectively. Colors distinguish the six flow velocities $v_{0}$ imposed by a thermostatted pump in the pore entrance (yellow shaded regions).
For Hagen–Poiseuille flow, the curvature of the velocity profile $v_{x}(z)$ equals the ratio of the pressure gradient $\Delta P/L$ and the shear viscosity $\eta$, and it does not depend on the boundary condition (i.e., the slip length ${\delta_{\text{slip}}}$); this is evident from the the velocity profiles in Eqs. 4 and 6. In order to test this relation in the NEMD simulations, we determined the fluid pressure $P(x)$ along the direction of the flow Fig. 6a)-b). The pressure exhibits an approximately linear decrease due to the viscous resistance and the surface stresses imposed on the fluid as it moves through the pore. For the pore with crystalline walls Fig. 6a), the pressure drop is less pronounced than in the case of amorphous walls Fig. 6b), which is inline with the picture of a lower resistance of the surfaces to the flow in the crystalline case. The slope of the pressure profiles increases (in magnitude) as the imposed flow velocity is increased, suggesting that the stronger driving of the fluid leads to enhanced friction forces and also higher energy dissipation in the fluid. A constant pressure gradient, $\partial_{x}P(x)\approx\Delta P/L$, is expected from the linearized Navier–Stokes equation for laminar and incompressible, unidirectional flows in the absence of an external body force, see Eq. 1. We attribute the obtained small deviations from a linear behaviour of $P(x)$ to the non-zero compressibility of the simulated fluid. Linear fits to the pressure data in Fig. 6 provide us with values for $\Delta P/L$ for each pumping velocity $v_{0}$ studied.
Figure 7: a) Effective shear viscosity $\eta$ obtained from fits to the Hagen–Poiseuille flow in Fig. 4 as a function of the pumping velocity $v_{0}$ for three pore widths: $D=15\sigma$ (disks), $D=30\sigma$ (diamonds), and $D=60\sigma$ (stars). Colors distinguish results for pores with the crystalline, red, and amorphous, blue, wall configurations. Lines connect data points of the same pore to guide the eye.  b) Effective shear viscosity $\eta$ at a flow velocity of $v_{0}=0.5/\sigma\tau^{-1}$ as function of the pore width $D$ for the two surface morphologies. The black line represents the shear viscosity $\eta_{b}$ of the bulk fluid, Eq. 36.
The simulated flows may be used to define an effective shear viscosity $\eta$ of the fluid as the ratio of $\Delta P/L$ and the curvature $\partial_{x}^{2}v_{x}(z)/2$ entering the parabolic fits to the velocity profiles in Fig. 4. The viscosities obtained for $D=15\sigma$ depend only mildly on the flow velocity $v_{0}$, with an exception at the smallest value of $v_{0}$, Fig. 5b). The values for the crystalline walls are slightly below those for amorphous walls; however, the overall magnitude $\eta\approx 0.5\varepsilon\tau\sigma^{-3}$ is smaller by almost a factor of 3 than the bulk value obtained in separate simulations of the homogeneous fluid, Eq. 36. In order to clarify this discrepancy, we have performed NEMD simulations for fluid flow inside much wider slit-shaped pores, using $D=30\sigma$ and $60\sigma$, and have again determined the effective viscosity for a range of pumping velocities, Fig. 7a). Again, there is the dependency on the surface morphology and some variability of $\eta$ as a function of $v_{0}$. More significantly, we find that $\eta$ increases systematically with the pore width $D$ such that, for $D=60\sigma$, the effective viscosity of the confined flow reaches about 80% of $\eta_{b}$, Fig. 7b).
IV.4 Permeability of the pore
Figure 8: a) Relation between the mass flux $J_{\varrho}$ imposed on the simulated fluid and the obtained pressure gradient $\Delta P/L$. Colors distinguish the two surface morphologies studied: crystalline walls, red, and amorphous walls, blue. Symbols encode the pore widths: $15\sigma$ (circles), $30\sigma$ (diamonds), and $60\sigma$ (stars). Straight solid lines are fits to the data, testing Darcy’s law, Eq. 10.  b) Permeability $k$ of the nanopore as a function of the pore width $D$. Symbols show the values obtained from the linear fits in panel (a) for amorphous (blue disks) and crystalline (red disks) walls for each pore width. Solid lines are fits of Eq. 13 to the data points with ${\delta_{\text{slip}}}=5.88\sigma$ and ${\delta_{\text{slip}}}=0.20\sigma$, for the crystalline and amorphous walls, respectively. The shaded regions around the data points reflect the uncertainty in the effective viscosity values (Fig. 7a), entering the calculation of $k$. The black dashed line represents the permeability, $k\propto D^{2}$, predicted for Hagen-Poiseuille flow in a slit-shaped conduit with no-slip boundary conditions, Eq. 11.
The mass flux $J_{\varrho}$ as a function of the pressure gradient is shown in Fig. 8 for the three different pore widths analysed, $D/\sigma=15,30,60$. The data exhibit a linear relationship for small pressure gradients, in agreement with Darcy’s law, Eq. 10. Nonlinear behavior, i.e., a deviation from Darcy’s law, begins to emerge at the largest pressure gradients obtained, particularly for the pore with $D=60\sigma$ and amorphous walls. In general, the mass flux for the crystalline pore walls, red symbols, is consistently higher than for amorphous walls, blue symbols, across all pressure gradients, indicating a higher permeability. This observation aligns with our findings on the slip length, Fig. 5b): In Hagen-Poiseuille flow through a slip-shaped conduit, the permeability $k=D^{2}/12$ is expected to increase quadratically with the width, Eq. 11. Allowing for partial slip reduces the friction at the surfaces and leads to an increase of the permeability, depending on the slip length ${\delta_{\text{slip}}}$, Eq. 13. Thus, the higher permeability observed for the crystalline walls is routed in the larger slip lengths compared to the amorphous walls. We note that the permeabilities for the widest pores studied ($D=60\sigma$) deviate considerably from the fluid mechanical predictions, even if partial slip is included.
IV.5 Heating due to viscous dissipation
The profiles of the fluid temperature $T(x)$ along the direction of the flow (Fig. 6c,d) provide insights into the energy dissipation of the fluid as it is pushed through the nanopore. For both crystalline and amorphous walls, the temperature increases approximately linearly with the flow, which we will justify below. For the smaller pumping velocities ($v_{0}\lesssim 0.25\sigma\tau^{-1}$), the fluid is near global thermal equilibrium and the slight increase in temperature is consistent with a moderate energy dissipation. However, at higher flow velocities ($v\gtrsim 0.5\sigma\tau^{-1}$), a pronounced rise of the temperature is observed, up to factors of 2 and 3, indicating a significant increase in heat generation, which we attribute to the higher frictional forces at play. The temperature rise is weaker by about 30% for the crystalline walls (Fig. 6c) than for the amorphous ones (Fig. 6d), which qualitatively aligns again with the increased surface slip in the first case.
The almost linear variation of pressure $P(x)$ and temperature $T(x)$ along the flow has consequences for the fluid density $\varrho(x)$ since, anticipating local thermal equilibrium, the three quantities are coupled via the thermal equation of state. The data exhibit also a linear change of the density for all simulated flows and wall configurations, Fig. 6e)-f); however, the overall variation of $\varrho(x)$ is comparably small, on the order of 15–25% for the largest flow velocity considered.
Finally, we provide an argument from fluid mechanics to support the linear increase of temperature. We recall that the NEMD simulations are performed such that the total energy is locally conserved in the pore, so that the generated heat needs to be transported downstream to the thermostatted pump. In particular, the fluid flow is isolated energetically and a heat transfer into the pore walls is not included in the simulated set-up.
The heat introduced into the fluid due to viscous stresses as it moves non-uniformly is given by the viscous dissipation function, Eq. 20, which reduces to $\Phi=\eta(\partial_{z}v_{x})^{2}$ for uni-directional, incompressible Newtonian flow. It is independent of the boundary conditions at the surfaces and reads explicitly for Hagen–Poiseuille flow, Eq. 4:
$\Phi(z)=\frac{(\partial_{x}P)^{2}}{\eta}\,z^{2}\,.$ (38)
For simplicity, we will treat the viscosity as being uniform in space, i.e., we neglect the dependencies of $\eta$ on temperature, density, and shear rate; noting that a more advanced, even non-local theory may be more realistic [104].
The dissipation implies a change in temperature that is determined by the energy equation, Eq. 28, which holds rather generally for laminar flow.8 Focusing on steady flow in slit geometry, it reads:
$c_{v}m\varrho(x,z)v_{x}(x,z)\partial_{x}T(x,z)\\ =\kappa(\partial_{x}^{2}T(x,z)+\partial_{z}^{2}T(x,z))+\Phi(z),$ (39)
To make further progress, we assume that the temperature is transversally equilibrated, i.e., in the pore cross section, so that we can approximate $\partial_{z}T(x,z)\approx\text{const}$ and set $\bar{T}(x):=T(x,z)$. In addition, we introduce the transversally averaged viscous dissipation:
[…]
In addition, we note that the mass flux $J_{\varrho}$ through the cross section of the pore [Eq. 8] is constant along a steady flow due to mass conservation [Eq. 18]; $J_{\varrho}$ appears on the l.h.s. of Eq. 39 after integration over $z$. This leads to a first-order ordinary differential equation in $\partial_{x}\bar{T}(x)$:
$\kappa\partial_{x}^{2}\bar{T}(x)-c_{v}J_{\varrho}\partial_{x}\bar{T}(x)=-\bar{\Phi}.$ (41)
Neglecting heat conduction, $\kappa\approx 0$ implies a constant temperature gradient:
$\partial_{x}\bar{T}=\frac{\bar{\Phi}}{c_{v}J_{\varrho}}\,,$ (42)
which in the case of Hagen–Poiseuille flow, upon inserting Darcy’s law and the permeability, Eqs. 10 and 11, simplifies to:
$\partial_{x}\bar{T}=\frac{\partial_{x}P}{m\varrho c_{v}}.$ (43)
Correspondingly, the temperature $\bar{T}(x)$ increases linearly.
V Summary and Conclusions
In summary, we have investigated the role that the atomic-scale morphology of the pore surfaces plays in the flow behavior of a simple fluid confined to slit-shaped nanopores. Utilizing boundary-driven NEMD simulations with large particle numbers, we examined two kinds of pore walls realized by crystalline and amorphous obstacle configurations, respectively; in the latter case, the wall structure was chosen to match that of the equilibrium fluid. We have shown that the two surface morphologies lead to very different surface resistances to fluid flow: whereas a pronounced, well-known slip occurs in the flow near the crystalline walls, high surface stresses near the amorphous walls drastically reduce the slip and lead to almost ideal no-slip boundary conditions (Fig. 4). Accordingly, the obtained permeabilities of pores with crystalline surfaces are constantly higher than in the case of amorphous surfaces (Fig. 8).
The different slip behavior correlates with the fluid structure near the surfaces as follows: near the crystalline walls, the fluid structure shows strong density variations and layering (Fig. 3a). However, at comparable distances to the amorphous walls, the fluid remains homogeneous, which is accompanied by an apparent depletion zone at the fluid–surface interface with a monotone decay of the density to zero (Fig. 3b). For both wall types, we observed that the fluid structure is independent of the flow velocity and resembles the equilibrium behavior, also at the fastest flows studied. This suggests that there is a unique effective surface potential seen by the fluid, which can be altered significantly by changing the surface morphology; in particular, attractive effective forces may appear despite the absence of attraction between the fluid and wall particles.
The simulated transversal velocity profiles $v_{x}(z)$, in the cross section of the pore, follow the predictions of Hagen-Poiseuille theory for incompressible, Newtonian fluids, provided that a slip length is included, Eq. 4. The obtained slip lengths ${\delta_{\text{slip}}}$ are almost independent of the flow velocity and characterize the pore surface (Fig. 5). However, the curvature of the parabolic velocity profiles suggests an effective shear viscosity $\eta$ that is significantly smaller than its bulk value, $\eta_{b}$. We attribute this effect to the non-zero compressibility of the simulated fluid and, second, to the nanoconfinement. Our data provide evidence for a systematic increase of $\eta(D)$ towards $\eta_{b}$ as the pore width $D$ is increased (Fig. 7b).
The fluid temperature increases linearly along the flow in our energy-conserving NEMD simulations (Fig. 6c,d). Such behavior may be explained by the dominant role of viscous dissipation, quantified by $\bar{\Phi}$, in combination with fluid convection, given by the mass flux $J_{\varrho}$. Within the fluid mechanics of Newtonian fluids, Eq. 39, the viscous dissipation is treated as a localized heat source that increases the fluid temperature proportionally to the viscous stresses. This leads to the prediction of a constant temperature gradient: $\partial_{x}\bar{T}\propto\bar{\Phi}/J_{\varrho}$, see Eq. 42. This line of argument is consistent with the data for $T(x)$, showing smaller thermal gradients for the crystalline pore walls as compared to the amorphous walls. The latter case implies a higher energy dissipation, which is attributed to the irregularity of the wall structure; for crystalline walls, on the other hand, the comparably flat surface minimizes the frictional resistance to flow. We acknowledge that an unbounded increase of the fluid temperature in the pore challenges physical intuition and seems to contradict experimental evidence. Yet, we note that (i) measurements of the fluid temperature within a nanopore are not available to our knowledge, and (ii) heat transfer to and heat conduction within the pore material is not included in the investigated, oversimplified simulation model. Such heat transfer obviously delimits the rise in temperature, depending on a set of additional parameters that one would need to introduce into the model along with additional non-trivial physical processes such as Kapitza heat resistance [105, 106, 107].
The fluid depletion near the amorphous surface (Fig. 3b) may suggest that such surfaces are solvophobic, which may appear counterintuitive given the almost ideal no-slip conditions are observed at this surface. However, it was pointed out previously [56] that there is no simple correlation between the wetting properties and the surface potential (and thus the fluid structure). Similarly, we have provided a simple example for the absence of surface slip that is not intuitively inferred from an effective surface potential. It would be interesting to test a proposed relation of the surface slip with the fluid structure and its transport coefficients near the surface [108]. Along these lines, the further characterization of amorphous, fluid-matched wall structures, is left for future research.
Our findings point on the importance of molecular-level details in designing nanoporous materials for specific fluid transport applications. Tailoring the surface interactions and their morphology of nanoporous materials has the potential to optimize the efficiency of nanoscale filtration, drug delivery, or energy storage.
Data availability statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Acknowledgements.
Financial support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 523950429 is gratefully acknowledged.